OpenStudy (anonymous):

Need Help What is (36 + 64)² = 36² + 2(36)(64) + 64²? Is the answer 5,000?

OpenStudy (anonymous):

Is anyone going to be able to help with this last problem

OpenStudy (anonymous):

add 36 and 64 then take squ

OpenStudy (anonymous):

I don't get where you get your 5000 from actually... On another note, it is true that: (36 + 64)² = 36² + 2(36)(64) + 64² or if you can think of it this way: let x=36, y=64 then: (x + y)² = x² + 2(x)(y) + y²

OpenStudy (anonymous):

they are not variables they are constants (100)^2 =10000

OpenStudy (anonymous):

I understand that, but there does not exist a question, though there is a suggested answer of 5000 from the info given, which is (36 + 64)² = 36² + 2(36)(64) + 64². which is 10000=10000....I fail to see the question in that unless it is to verify the relation I have stated on my previous post.